# Evaluation of Language Models in the Medical Context Under Resource-Constrained Settings

### v1

@@ 4 Methodology
The experiments involve two target tasks: text classification and conditional text generation. Specifically, text classification is addressed through three different approaches: (i) contextual embedding similarity, (ii) natural language inference (NLI), and (iii) multiple-choice question-answering (MCQA). These are explained in Section 4.1.2. As detailed below, each task is evaluated using diverse datasets, models, and performance metrics.
The evaluation process comprises 53 models, which exhibit a size spectrum ranging from 110 million to 13 billion parameters. These models belong to three knowledge domains: general, biomedical or scientific, and clinical. Their characterization can be found in Table 1. OpenLLaMA 13B, Flan-T5-XXL, T5-V1.1-11B, and T0++ are run using float16 precision.
In order to conduct a comprehensive analysis, we recognize the importance of evaluating the chosen models across a broad spectrum of medical specialties, ultimately aiming at a general assessment of medical knowledge. To achieve this, we used the transcriptions dataset. Nevertheless, also recognizing the potential of artificial intelligence in particular medical fields, we narrowed our focus in addition to radiology, one of the most promising fields for AI integration. Consequently, we opt for the MIMIC-CXR dataset along with its labeled counterpart, MS-CXR.
| Small (S) | Medium (M) | Large (L) | XL | XXL | ||||||||||||
| ID | Model | Size | ID | Model | Size | ID | Model | Size | ID | Model | Size | ID | Model | Size | ||
| Encoder-only | General | m00 | BERT 0 [14] | 110 M | m01 | BERT 0 [14] | 340 M | - | - | - | - | - | - | - | - | - |
| Biomedical | m02 | BiomedBERT | 110 M | m04 | BiomedBERT-large | 340 M | - | - | - | - | - | - | - | - | - | |
| (abstracts + full text) 0 [71] | (abstracts only) 0 [71] | |||||||||||||||
| m03 | BiomedBERT | 110 M | m12 | RoBERTa-MNLI ‡1 [81] | 355 M | |||||||||||
| (abstracts only) 0 [71] | ||||||||||||||||
| m05 | SciBERT 0 [70] | 110 M | ||||||||||||||
| m06 | SapBERT 0 [82] | 110 M | ||||||||||||||
| m07 | BioLORD-STAMB2-v1 0 [83] | 110 M | ||||||||||||||
| m08 | BioLORD-STAMB2-v1-STS2 0 [83] | 110 M | ||||||||||||||
| m09 | BioLORD-PMB 0 [83] | 110 M | ||||||||||||||
| m11 | NLI-DeBERTa ‡1 [84] | 100 M | ||||||||||||||
| Clinical | m10 | Bio+Clinical BERT 0 [75] | 110 M | - | - | - | - | - | - | - | - | - | - | - | - | |
@@
decoder |
General | m14 | T5-V1.1-Base 2 [85, 86] | 220 M | m13 | BART Large-MNLI 1 [87] | 407 M | m15 | T5-V1.1-Large 2 [85, 86] | 770 M | m16 | T5-V1.1-3B 2 [85, 86] | 3.0 B | m17 | T5-V1.1-11B 2 [85, 86] | 11.0 B |
| m18 | Flan-T5-Base †2 [15] | 220 M | m19 | Flan-T5-Large †2 [15] | 770 M | m20 | Flan-T5-XL †2 [15] | 3.0 B | m21 | Flan-T5-XLL †2 [15] | 11.0 B | |||||
| m22 | T0 3B †2 [88] | 3.0 B | m23 | T0++ †2 [88] | 11.0 B | |||||||||||
| Biomedical | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
| Clinical | m24 | ClinicalT5-base 2 [89] | 220 M | - | - | - | m25 | ClinicalT5-large 2 [89] | 700 M | - | - | - | - | - | - | |
| Decoder-only | General | - | - | - | m26 | GPT-2 Medium3 [90] | 355 M | m27 | GPT-2 Large 3 [90] | 774 M | m28 | GPT-2 XL 3 [90] | 1.5 B | m29 | Palmyra Base 5B 23 [91] | 5.0 B |
| m41 | OpenLLaMA 3B 3 [92] | 3.0 B | m30 | Camel 5B†2 [93] | 5.0 B | |||||||||||
| m42 | OpenLLaMA 3Bv2 3 [92] | 3.0 B | m31 | GPT-J 6B 23 [94] | 6.0 B | |||||||||||
| m32 | Instruct GPT-J †2 [95] | 6.0 B | ||||||||||||||
| m33 | Falcon-7B 23 [96] | 7.0 B | ||||||||||||||
| m34 | Falcon-7B-Instruct †2 [96] | 7.0 B | ||||||||||||||
| m35 | MPT-7B 23 [97] | 7.0 B | ||||||||||||||
| m36 | MPT-7B-Instruct †2 [97] | 7.0 B | ||||||||||||||
| m37 | LLaMA-7B 23 [25] | 7.0 B | ||||||||||||||
| m38 | LLaMA 2-7B 23 [26] | 7.0 B | ||||||||||||||
| m39 | Alpaca 7B †2 [98] | 7.0 B | ||||||||||||||
| m40 | LLaMA 2-CHAT-7B †2 [26] | 7.0 B | ||||||||||||||
| m43 | OpenLLaMA 7B 3 [92] | 7.0 B | ||||||||||||||
| m44 | OpenLLaMA 7Bv2 3 [92] | 7.0 B | ||||||||||||||
| m45 | OpenLLaMA 13B 3 [92] | 13.0 B | ||||||||||||||
@@
/ Scientific |
- | - | - | m48 | BioGPT 3 [74] | 347 M | m47 | GPT-2-PubMed Large 3 [99] | 774 M | m50 | Galactica 1.3B 3 [24] | 1.3 B | m51 | Galactica 6.7B 3 [24] | 6.7 B | |
| m46 | GPT-2-PubMed Medium 3 [99] | 355 M | m49 | BioGPT-Large 3 [74] | 1.5 B | |||||||||||
| Clinical | - | - | - | - | - | - | - | - | - | - | - | - | m52 | MedAlpaca 7b †2 [58] | 7.0 B | |
@@ 4.1 Text classification
For the text classification task, we evaluate two datasets using three approaches and their corresponding models. It is essential to mention that fine-tuning is not performed; however, we use zero-shot prompt-engineering. In the following, we provide a concise overview of the datasets, followed by an explanation of the approaches employed. We refer the reader to Appendix A for more details on the preprocessing steps and the data characteristics of the datasets.
@@ 4.1.1 Datasets
Transcriptions is a multi-label collection of electronic health records covering several medical specialties. The preprocessing procedure for this dataset involves removing null entries, modifying the report’s format, and selecting and renaming labels. After preprocessing, the resultant dataset consists of 2074 samples and 29 labels. Model performance is evaluated using the AUC score.
In particular, since the EHRs are used in their entirety, certain samples contain token vectors that exceed the maximum input length allowed for multiple models. There are even samples whose token vectors exceed 2048 tokens, the maximum input length supported among the models. Consequently, different strategies have been adopted for each approach to effectively address this challenge. The adopted strategies, which require splitting the sentence into parts, will produce different results than processing the entire sample at once due to the attention mechanisms. These approaches are detailed in Section 4.1.2.
MS-CXR is a multi-class dataset comprising sections of X-ray reports, each accompanied by annotations made by a radiologist [100, 101, 102]. The number of classes is eight, well distributed and represented in 718 unique text samples. The preprocessing procedure is more straightforward than for the other datasets and includes removing samples with missing information and duplicates. In contrast to transcriptions, there is no evidence of any sample exceeding the maximum input length allowed for any of the models.
Model performance is evaluated using a set of metrics in their macro-averaged version to ensure a comprehensive assessment. The metrics are accuracy, F1-score, precision, and recall.
@@ 4.1.2 Approaches
Contextual embedding similarity is grounded in the cosine similarity between the contextual embeddings of the text under consideration and its corresponding label. The models employed are exclusively encoder-only, comprising 11 models subjected to evaluation. The final contextual or sentence embedding is determined using three distinct pooling strategies: the CLS-token embedding, average pooling of the token-level embedding, and maximum pooling of the token-level embedding. In addition, prompt-engineering is implemented for the label. The specific details of the proposed prompts can be found in Appendix B.
In cases where the input tokens exceed the maximum input size defined by the model, the samples are split into smaller parts, processing each individually. The embedding of the sample varies depending on the pooling strategy implemented, as described below. In particular, the models used in this approach have a maximum input token size of 512 tokens.
CLS pooling: The contextual embeddings of the constituent parts of a sample are obtained by CLS pooling. The contextual embedding of the sample is calculated as the element-wise average among the contextual embeddings of the constituent parts.
@@
Maximum pooling: The contextual embedding of each sample is determined as the element-wise maximum of the output token-level embeddings of the constituent parts.
Average pooling: The average operator considers the number of non-padding tokens of the parts into which a sample’s text is segmented. Thus, the contextual embedding of each sample is calculated as the element-wise weighted average among the token-level embeddings of its constitutive parts.
Natural language inference for text classification involves representing the text to be classified as a premise and the labels as hypotheses. Subsequently, a multi-classification task with three classes -entailment, neutral, and contradiction- is performed. Conceptually, this approach can be thought of as having classifiers, where corresponds to the number of labels or classes. The models employed are encoder-only (cross-encoders) or encoder-decoder architectures.
How the output is processed varies depending on whether the dataset is multi-label or multi-class. The entailment and contradiction outputs are transformed into binary probabilities in the multi-label scenario. This transformation effectively translates the problem into binary text classifiers, each classifier for one of the labels. In contrast, in the multi-class scenario, only the entailment logits are used to predict the class associated with a sample. When the length of the input tokens exceeds the maximum input size defined by the model, the scores of these samples are calculated by aggregating the scores obtained from their constituent parts. Aggregation measures that can be considered include maximum, average, or median.
Furthermore, like the previous approach, the influence of prompts is considered during the model evaluation. The prompts used in this context are identical to those employed in the contextual embedding similarity approach, and their details can be found in Appendix B.
@@
Multiple-choice question-answering framework enables encoder-decoder and decoder-only architectures to perform text classification. A total of 27 models are assessed in this approach. Table 1, which provides details of the models used, shows the inclusion of pre-trained models and their instruction-tuned counterparts.
For this approach to work, a prompt is required. Therefore, the impact of the prompt is measured in the analysis. The prompts are formulated based on those most commonly used in instruction-tuning models for multiple-choice question-answering tasks. However, in these prompts, the question format is not restrained. For this reason, this study also considers the impact of the question itself. Additional information on the prompts used in this approach can be found in Appendix B.
It is essential to mention that the logit space has been constrained to align with the set of viable alternatives presented in a multiple-choice scenario. For this, the token identifiers associated with the feasible response options are determined and used to filter the logit space.
4.2 Conditional text generation
@@
We evaluate one dataset for the conditional text generation task, using perplexity as the evaluation metric. Perplexity is a measure of uncertainty on the value of a sample from a discrete probability distribution. If we have a tokenized sequence , then
where is the log-likelihood of the -th token conditioned on the preceding tokens according to our model.
Below, a concise overview of the dataset employed is provided, followed by an explanation of the approach adopted. A more detailed characterization of MIMIC-CXR and its preprocessing procedure is presented in Appendix A.
4.2.1 Datasets
MIMIC-CXR is composed of X-ray reports [103, 104, 102]. As part of the preprocessing procedure, the relevant sections are extracted using the code provided by Johnson et al. [105, 106]. Following this extraction process and the subsequent removal of null and duplicated samples, the resulting dataset contains samples. Furthermore, none of the samples exceed the maximum input size allowed for the proposed models.
4.2.2 Approaches
@@
For this task, 20 decoder-only models are evaluated, as detailed in Table 1. Among these models, we find Galactica, whose tokenizer lacks special tokens. As a result, the analysis is conducted in two parts. In the first part, the Beginning of Sequence (BOS) token is included during tokenization, with perplexity calculated from the first token of the texts. The BOS token is a special token usually employed by generative models to indicate the start of a text. In cases where a model’s tokenizer does not have the BOS token predefined, such as the Falcon-7B, it is defined as its first special token. The second part of the analysis proceeds without the BOS token, and perplexity is determined from the second token of the texts, and so on.

### v2

@@ 4 Methodology
A series of experiments on medical text classification and conditional text generation are carried out to understand better the behavior of language models under resource-constrained settings, i.e., settings with consumer-grade computing resources. In total, language models are evaluated, whose size ranges from million to billion parameters. The selection of these models spans the general, biomedical, and clinical knowledge domains and includes the three families of Transformer-based models. Moreover, only open-source, smaller than B parameters models are considered. Details on the selected models are found in Table 1 and Appendix B.
All experiments are performed using a Quadro RTX 8000 GPU and CUDA version 12.2. To guarantee that the selected models align with consumer-grade computing resources, models with more than billion parameters (i.e., OpenLLaMA 13B, Flan-T5-XXL, T5-V1.1-11B, and T0++) are run with float precision. By halving the floating-point precision, these and billion parameter model versions are still viable in computational resource-constrained settings.
The three families of Transformer-based models are considered for the text classification task via different approaches (described in Section 4.1.2), whereas solely decoder-only models are used for the conditional text generation task. Transcriptions, MIMIC-CXR, and MS-CXR have been chosen as evaluation datasets. Transcriptions covers a broad spectrum of medical specialties, allowing a general assessment of medical knowledge. MIMIC-CXR and its labeled version, MS-CXR, enable testing focused on radiology, one of the most promising fields for AI integration, narrowing the evaluation to specialized medical knowledge.
Small (S)
Medium (M)
Large (L)
XL
XXL
ID
Model
Size
ID
Model
Size
ID
Model
Size
ID
Model
Size
ID
Model
Size
Encoder-only
General
m00
BERT 0 [16]
110 M
m01
BERT 0 [16]
340 M
m11
NLI-DeBERTa ‡1 [88]
100 M
m12
RoBERTa-MNLI ‡1 [89]
355 M
Biomedical
m02
BiomedBERT
110 M
m04
BiomedBERT-large
340 M
(abstracts + full text) 0 [66]
(abstracts only) 0 [66]
m03
BiomedBERT
110 M
(abstracts only) 0 [66]
m05
SciBERT 0 [65]
110 M
m06
SapBERT 0 [90]
110 M
m07
BioLORD-STAMB2-v1 0 [91]
110 M
m08
BioLORD-STAMB2-v1-STS2 0 [91]
110 M
m09
BioLORD-PMB 0 [91]
110 M
Clinical
m10
Bio+Clinical BERT 0 [70]
110 M
@@
decoder
General
m14
T5-V1.1-Base 2 [92, 93]
220 M
m13
BART Large-MNLI 1 [94]
407 M
m15
T5-V1.1-Large 2 [92, 93]
770 M
m16
T5-V1.1-3B 2 [92, 93]
3.0 B
m17
T5-V1.1-11B 2 [92, 93]
11.0 B
m18
Flan-T5-Base †2 [17]
220 M
m19
Flan-T5-Large †2 [17]
770 M
m20
Flan-T5-XL †2 [17]
3.0 B
m21
Flan-T5-XLL †2 [17]
11.0 B
m22
T0 3B †2 [95]
3.0 B
m23
T0++ †2 [95]
11.0 B
@@
Clinical
m24
ClinicalT5-base 2 [96]
220 M
m25
ClinicalT5-large 2 [96]
700 M
Decoder-only
General
m26
GPT-2 Medium3 [19]
355 M
m27
GPT-2 Large 3 [19]
774 M
m28
GPT-2 XL 3 [19]
1.5 B
m29
Palmyra Base 5B 23 [97]
5.0 B
m41
OpenLLaMA 3B 3 [98]
3.0 B
m30
Camel 5B†2 [99]
5.0 B
m42
OpenLLaMA 3Bv2 3 [98]
3.0 B
m31
GPT-J 6B 23 [100]
6.0 B
m32
Instruct GPT-J †2 [101]
6.0 B
m33
Falcon-7B 23 [102]
7.0 B
m34
Falcon-7B-Instruct †2 [102]
7.0 B
m35
MPT-7B 23 [103]
7.0 B
m36
MPT-7B-Instruct †2 [103]
7.0 B
m37
LLaMA-7B 23 [27]
7.0 B
m38
LLaMA 2-7B 23 [28]
7.0 B
m39
Alpaca 7B †2 [104]
7.0 B
m40
LLaMA 2-CHAT-7B †2 [28]
7.0 B
m43
OpenLLaMA 7B 3 [98]
7.0 B
m44
OpenLLaMA 7Bv2 3 [98]
7.0 B
m45
OpenLLaMA 13B 3 [98]
13.0 B
Biomedical
/ Scientific
m48
BioGPT 3 [69]
347 M
m47
GPT-2-PubMed Large 3 [105]
774 M
m50
Galactica 1.3B 3 [26]
1.3 B
m51
Galactica 6.7B 3 [26]
6.7 B
m46
GPT-2-PubMed Medium 3 [105]
355 M
m49
BioGPT-Large 3 [69]
1.5 B
Clinical
m52
MedAlpaca 7b †2 [71]
7.0 B
@@ 4.1 Text classification
Text classification is addressed using the Transcriptions and MS-CXR datasets and three different approaches: (i) contextual embedding similarity, (ii) natural language inference (NLI), and (iii) multiple-choice question answering (MCQA). The contextual embedding similarity approach is intended for encoder-only models, the NLI approach for encoder-only and encoder-decoder models pre-trained for NLI, and the MCQA approach for encoder-decoder and decoder-only models..
Model tuning is implemented through zero-shot learning. To analyze the impact of prompting on text classification performance, different prompts are applied during inference. These prompts, grouped into two sets, are defined according to the classification approach. The first set of prompts is used for contextual embedding similarity and NLI. Since neither of these approaches requires a prompt to work, its non-use is also included in the analysis. The second set of prompts is used for MCQA, approach that needs a prompt to work. Prompts from the second set are defined based on those most commonly used in instruction-tuning models for multiple-choice question answering tasks.
Let be a text sample and be a class, not necessarily corresponding to . A prompt from the first prompt set, , is defined as a function of a prompt template and a label. For example, “This is an example of ”. The set is only applied to the classes. Meanwhile, a prompt from the second prompt set, , combines a prompt template (consisting of the prompt structure and a question), a text sample, and the classes. For example, “You are a doctor and have the following information about a patient from a chest x-ray: . What is the diagnosis? . (”. In this example, the prompt template consists of the question “What is the diagnosis?” and the prompt structure, which is the rest of the text. Prompts are presented in detail in Appendix C.
@@ 4.1.1 Datasets
The datasets evaluated in text classification are Transcriptions and MS-CXR. Each of these datasets is introduced below. Their preprocessing and characterization details are given in Appendix A.
Transcriptions is a multi-label collection of electronic health records (EHRs) covering many medical specialties. Preprocessing is applied to the data, removing null entries, organizing the EHR format, and selecting the final set of labels. After all, samples and classes are available. Performance is measured by the AUC score since the dataset is multi-label.
Due to the length of some EHRs, certain token vectors exceed the maximum input length allowed by some models. To cope with this limit, the input sequence is processed using a non-overlapping sliding window method [87], as detailed in Section 4.1.2.
MS-CXR is a multi-class dataset composed of X-ray report sections, each accompanied by annotations made by a radiologist [106, 107, 108]. There are unique samples representing eight well-distributed classes. Preprocessing of this dataset consists of removing samples with missing information and duplicates. Contrary to Transcriptions, no sample exceeds the maximum allowed input length for any of the models. Performance is measured by accuracy, F1-score, precision, and recall in their macro-averaged version to ensure a comprehensive assessment.
@@ 4.1.2 Approaches
Text classification is performed through (i) contextual embedding similarity, (ii) natural language inference, and (iii) multiple-choice question answering.
@@
Contextual embedding similarity is grounded in the cosine similarity between the contextual embeddings of the sample text and the classes. The contextual or sentence embedding is determined by three distinct pooling strategies: CLS-token embedding, average token-level embedding pooling, and maximum token-level embedding pooling.
For this approach, encoder-only models are employed, with a total of models evaluated. These models have a maximum input token size of tokens. Therefore, the samples’ token vectors that exceed this limit are processed with the non-overlapping sliding window method. The fragments are aggregated according to the pooling strategy, as follows.
@@
CLS pooling: The contextual embedding of each fragment from a sample is computed as its CLS-token output embedding. These embeddings are then aggregated using the element-wise average to obtain the contextual embedding representing the sample.
Maximum pooling: The contextual embedding of each fragment from a sample is computed by applying element-wise maximum at token level over the output embeddings. These embeddings are then aggregated using again the element-wise maximum to obtain the contextual embedding representing the sample.
Average pooling: The contextual embedding of each fragment from a sample is computed by applying the element-wise average at the token level over the output embeddings. These embeddings are then aggregated using the element-wise weighted average to obtain the contextual embedding representing the sample. Average’s weights indicate the number of non-padding tokens in each fragment.
Natural language inference is the task of determining whether a hypothesis is true (entailment), false (contradiction), or indeterminate (neutral) given a premise. When applied for text classification, the premise represents a test sample, and the hypothesis represents the classes. For multi-class datasets, the predicted label is calculated from the entailment logits of each hypothesized class. For multi-label datasets, the entailment and contradiction logits are transformed into binary probabilities, which indicate whether or not a particular hypothesized class is predicted. This could be viewed as having binary text classifiers, where is the number of classes.
This approach employs encoder-only (cross-encoder) and encoder-decoder models. These models have a lower maximum input token size than some of the test samples. Therefore, the token vectors of these particular samples are processed with the non-overlapping sliding window method. They are divided into fragments, whose scores are calculated individually, and then these scores are averaged to get the score of the whole sample.
@@
Multiple-choice question answering enables generative models, i.e., encoder-decoder and decoder-only models, to perform text classification. A total of models are assessed in this approach, including both pre-trained models and their instruction-tuned versions.
As multiple-choice question answering is not intended for a extensive number of choices, the Transcriptions dataset is evaluated using a reduced version with eleven classes instead of the 29 available. These eleven classes consist of the ten most frequent labels plus an “Other” class. The number of samples evaluated is not affected. Additionally, the models’ logit space has been constrained to align with the response options of a multiple-choice scenario and, thereby, allow for automated evaluation. The token identifiers associated with the feasible response options are determined and used to filter the logit space.
4.2 Conditional text generation task
@@
Conditional text generation is assessed with the MIMIC-CXR dataset, using perplexity as the performance evaluation metric. Perplexity (PPL) is a measure of uncertainty on the value of a sample from a discrete probability distribution. Let be a tokenized sequence, then
where is the log-likelihood of the -th token conditioned on the preceding tokens .
@@
Decoder-only models are employed for evaluation, with a total of models considered. Among these models is Galactica, whose tokenizer lacks special tokens. As a consequence, two scenarios are analyzed: the inclusion and the non-inclusion of the start-of-sequence (BOS) token. The BOS token is a special token typically used by generative models to indicate the start of a text. In the first scenario, this token is included during tokenization, and perplexity is calculated from the first token in the texts. When the model’s tokenizer does not have the BOS token predefined, such as Falcon-7B, it is then defined as the tokenizer’s first special token. In the second scenario, the BOS token is excluded, and perplexity is calculated from the text’s second token.
4.2.1 Dataset
The dataset evaluated in conditional text generation is MIMIC-CXR, introduced below. Details on its preprocessing and characterization are in Appendix A.
MIMIC-CXR is an X-ray reports dataset [109, 110, 108]. Relevant sections of these reports are extracted using the code provided by Johnson et al. [111, 112]. Subsequently, null and duplicate samples are removed, with the resulting dataset having samples. None of these samples exceeds the maximum input size allowed for the proposed models.